2018-2019 Winter FiSV 2000 Mid-Term Exam- Par: 1 indlivicual Cha. ea Datto Draw Design Layout References Mailings Review View Help Tell me what you want to do Find . AaBbCeD AaBbCcDd AaBbCcDd AaBbCcD Replace 1仨.1c2.H. | | Heading i 1Normal 1No Spac.. Heading 2 | Select Par 反 Quick Ratio: -42 2. Shelton, Inc., has sales of $17.5 million, total assets of $13.1 million, and total debt of $5.7 million. If the profit margin is 6 percent, what is net income? What is ROA? What is ROE? Calculations: Answers: Net Income: ROA: ROE: -- 5 名
Answer #1

Net income = Total sales * Profit margin = $17.5 mn * 6% = $1.05 million

Net equity = Total assets - debt = 13.1 - 5.7 = $7.4 mn

ROA = NET INCOME/TOTAL ASSETS = 1.05/13.1 = 0.08015 or 8.015%

ROI = Net income/Net equity = 1.05/7.4 = 0.14189 or 14.189%